OpenStudy (anonymous):

Which of the following is the simplified form ofthe seventh root of x times the seventh root of x times the seventh root of x ?

OpenStudy (anonymous):

I thought my answer made the most sense since when the denominators are the same you keep them the same, right? then there was no numerator to go off of.

OpenStudy (kenshin):

remember xth root is just power of (1/x), and timesing by itself means to the power of 2, so timesing twice is a power of 3

OpenStudy (anonymous):

I am confused... haha

OpenStudy (kenshin):

a^(1/x) * a^ (1/x) = a^(2/x) not a^(2/2x)

OpenStudy (anonymous):

I think the answer should have been A. I am looking back on my notes now.

OpenStudy (kenshin):

exponents add when you multiply the bases and minus when you divide them.

OpenStudy (anonymous):

so since 7 is the base i add?

OpenStudy (kenshin):

so 7^(1/3) * 7^(1/3) should be 7^(1/3 + 1/3), you add the powers not the 7s

OpenStudy (kenshin):

and you can only do that when the bases are the same.

OpenStudy (anonymous):

wait, i am just simplifying so i shouldn't even put it into a radical expression

OpenStudy (kenshin):

yes, but to help see better, do you understand how seventh root of x = x^(1/7)?

OpenStudy (anonymous):

it's hard for me to read the equation

OpenStudy (kenshin):

$$x^{1/7}$$

OpenStudy (kenshin):

is that better, this is the way to express seventh root of x, are you familiar with this?

OpenStudy (anonymous):

yes

OpenStudy (kenshin):

ok then as I was saying, when you have $$x^{1/7}$$ times $$x^{1/7}$$, you do $$x^{1/7 + 1/7}$$, do you understand this step?

OpenStudy (anonymous):

i do

OpenStudy (kenshin):

then timesing 3 times will be adding their powers 3 times, and that would be?

OpenStudy (anonymous):

3/7? going out on a limb here cause i'm lost :/

OpenStudy (kenshin):

you're right

OpenStudy (anonymous):

surprising.

OpenStudy (anonymous):

x 3/7

OpenStudy (anonymous):

thank you much for your help, just trying to understand this stuff